[figure: 9.7 cm]

UNUSUAL FEATURES OF ITINERANT ANTIFERROMAGNETISM IN UGA_{3}

D. Kaczorowski*

Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P. O. Box 1410, 50-950 Wrocław, Poland

Antiferromagnetic UGa_{3} is a prototypical representative of an unusual class of itinerant magnets with large orbital moments. Although the magnetic, electrical transport and thermodynamic properties of this compound closely resemble those of 3d-electron transition-metal systems, some peculiar features unambiguously manifest the presence of large orbital moments, as in lanthanide-based materials with localized 4f-electrons. This unique behavior of itinerant 5f-electrons makes UGa_{3} an interesting system for both experimental and theoretical studies. In the talk, we shall review the results of a series of sophisticated experiments performed on high-quality single crystals of UGa_{3}: elastic and inelastic neutron scattering, optical ellipsometry, resonant X-ray magnetic scattering, Mössbauer spectroscopy. Among other topics we shall discuss a two-component dynamical magnetic response, related to fast and slow quasiparticle excitations in the ordered state, and a giant enhancement of the RXMS signal at the K-edge of nonmagnetic Ga anion, which possibly involves the hybridization of the gallium 4p-orbitals with the uranium 5f-states. The experimental findings will be compared to the calculated results, derived in the framework of local spin density functional theory. *Work done in collaboration with M. M. Abd-Elmeguid, N. Bernhoeft, M. S. S. Brooks, A. Czopnik, A. Hiess, G. H. Lander, P. M. Oppeneer, J. P. Sanchez, J. Schoenes.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : invited

Corresponding author : D. Kaczorowski

Address for correspondence : Institute of Low Temperatures and Structure Research, Polish Academy of Sciences, P. O. Box 1410, 50-950 Wrocław, Poland

Email address : D.Kaczorowski@int.pan.wroc.pl